Using y-mx+b write an equation for a line that goes through the poings (8,0) (6, -8).
i would say that m=0-(-8)/8-6 which is 8/2 or 4
so -8=4(0)+m

Answer by ewatrrr(24785)

 
Hi,
equation for a line that goes through the points:
(8, 0)
(6,-8) m = 8 / 2 = 4
y = 4x + b
0 = 4*8 + b
b = -32
y = 4x - 32